# From graph to Riesz continuity

### v2

@@ 2 Preliminaries
In fact, this reasoning provides a stronger result. It concerns another natural topology on $\mathcal{R}(H,H^{\prime})$, which is is strictly coarser then the graph topology, induced by the inclusion $\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow{{}^{s}{\mathcal{P}}}(H\oplus H^{\prime})$ from the strong operator topology on the space of projections. On the subset of self-adjoint operators it coincides with the strong resolvent topology, which is induced by the map $A\mapsto(A+i)^{-1}$, or equivalently by the map $A\mapsto(A-i)^{-1}$, from the strong operator topology on bounded operators. The reasoning in the proof of Proposition 2.1 shows that the bounded transform $\bm{\chi}\colon\mathcal{R}\to\mathcal{D}$ is continuous with respect to this coarser topology on $\mathcal{R}$ and the strong* topology on $\mathcal{D}$.
Proposition 2.1 follows from a similar result for self-adjoint operators and the restriction of $\bm{\chi}$ to $\mathcal{R}^{\mathrm{sa}}$, $A\mapsto A(1+A^{2})^{-1/2}$, by the standard trick of passing from $A$ to the self-adjoint operator $\left(\begin{smallmatrix}0&A^{*}\\ A&0\end{smallmatrix}\right)$. For self-adjoint operators, a similar reasoning provides continuity of an arbitrary function $f\in C_{b}(\mathbb{R})$ with respect to the strong resolvent topology on $\mathcal{R}^{\mathrm{sa}}$ and the strong topology on $\mathcal{D}$. A different proof of this fact based on the spectral theorem is given by [RS], Theorem VIII.20(b). A similar result for compact spaces of parameters is given by [W], Proposition 1.1 of Ch. Wahl.
Action of the unitary groups.   The unitary groups $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ act on $\mathcal{R}(H,H^{\prime})$ by the right and left multiplication respectively. These actions preserve the subsets $\mathcal{R}_{F}$ and $\mathcal{R}_{K}$. For $u\in\mathcal{U}(H^{\prime})$ and $v\in\mathcal{U}(H)$ we have
@@
so the actions of $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ (equipped with the norm topology) are continuous on both ${{}^{r}\mathcal{R}}$ and ${{}^{g}\mathcal{R}}$.
Recall that ${{}^{s}{\mathcal{U}}}(H)$, the unitary group equipped with the strong operator topology, is a topological group. Indeed, the multiplication map $\mathcal{B}(H)\times\mathcal{B}(H)\to\mathcal{B}(H)$, $(a,b)\mapsto ab$, is strongly continuous on $S\times\mathcal{B}(H)$ for every bounded subset $S$ of $\mathcal{B}(H)$. In particular, it is strongly continuous on $\mathcal{U}(H)\times\mathcal{U}(H)$. For every $u,u_{0}\in\mathcal{U}(H)$ and $\xi\in H$,
$\left\|(u^{-1}-u_{0}^{-1})\xi\right\|=\left\|u(u^{-1}-u_{0}^{-1})\xi\right\|=\left\|(u-u_{0})u_{0}^{-1}\xi\right\|,$
so the inverse map $\mathcal{U}(H)\to\mathcal{U}(H)$, $u\mapsto u^{-1}=u^{*}$, is also strongly continuous.
The actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{r}\mathcal{R}}$ and ${{}^{g}\mathcal{R}}$ are not continuous. However, continuity of their actions on $\mathcal{K}(H\oplus H^{\prime})$ provides the following result.
@@
The actions of the groups ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$ are continuous
@@
(2.7) $\operatorname{Gr}_{K}(H\oplus H^{\prime})\cong\mathcal{P}_{K}(H\oplus H^{\prime})=\left\{p\in\mathcal{P}(H\oplus H^{\prime})\,|\,p-p_{\infty}\text{ is compact}\right\},$
@@
3 Non-self-adjoint operators
[…]
we obtain a desired map $v\colon{{}^{g}\mathcal{R}}\to{{}^{s}{\mathcal{U}}}(H)$. If $\Gamma$ is considered as a Hilbert bundle over ${{}^{r}\mathcal{R}}$, then $\mathbf{w}$ becomes a trivialization in the category of bundles with the structure group $\mathcal{U}(H)$, so $v$ is continuous as a map ${{}^{r}\mathcal{R}}\to\mathcal{U}(H)$ as well.
@@ 3 Non-self-adjoint operators
=$\mathcal{I}=\left\{u\in\mathcal{B}(H,H\oplus H^{\prime})\,|\,u^{*}u=\operatorname{Id}_{H}\right\}$
@@
$\bm{\pi}\colon\mathcal{I}\to\operatorname{Gr}(H\oplus H^{\prime}),\quad u\mapsto\operatorname{Ran}(u).$
takes an isometry $u$ to its range; the corresponding projection is $uu^{*}$. The image $\bm{\pi}(\mathcal{I})$ of this map consists of infinite-dimensional subspaces of $H\oplus H^{\prime}$. We will need the following simple corollary of Kuiper’s theorem [Ku] about contractibility of $\mathcal{U}(H)$.
Lemma 3.1.
The map $\bm{\pi}\colon\mathcal{I}\to\bm{\pi}(\mathcal{I})$ is a principal $\mathcal{U}(H)$-bundle admitting a continuous section (and thus is trivial).
Proof. The group $\mathcal{U}(H)$ acts on $\mathcal{I}$ by the right multiplication. Every two isometries with the same range differ by an uniquely defined unitary operator $H\to H$, so this action turns $\bm{\pi}$ into a principal $\mathcal{U}(H)$-bundle. This bundle has a section by the same standard arguments as were used in the beginning of Section 2. Indeed, for every $u_{0}\in\mathcal{I}$ and $L_{0}=\bm{\pi}(u_{0})$ there is a continuous map $g$ from a neighborhood $B$ of $L_{0}$ to $\mathcal{U}(H\oplus H^{\prime})$ such that $L=g_{L}L_{0}$ for $L\in B$. The rule $L\mapsto g_{L}\cdot u_{0}$ determines a local section of $\bm{\pi}$ over $B$. Hence $\bm{\pi}$ is a locally trivial principal bundle over $\bm{\pi}(\mathcal{I})$ with the contractible structure group $\mathcal{U}(H)$. The base space $\bm{\pi}(\mathcal{I})$ is metric and thus paracompact. Therefore, $\bm{\pi}$ has a global section. $\square$
Proof of Theorem A.   The graphs of regular operators have infinite dimension, so ${{}^{g}\mathcal{R}}\subset\bm{\pi}(\mathcal{I})\subset\operatorname{Gr}(H\oplus H^{\prime})$. By Lemma 3.1, the restriction of $\bm{\pi}$ to ${{}^{g}\mathcal{R}}$ has a section. For the rest of the proof we fix such a section $u\colon{{}^{g}\mathcal{R}}\to\mathcal{I}$. In other words, $u$ is a graph-to-norm continuous map taking a regular operator $A$ to an isometry $u_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$.
@@
Remark.
Our construction of $v$ depends on the choice of a section $u$. Since $\bm{\pi}$ is a principal $\mathcal{U}(H)$-bundle, any other choice of a section differs from $u$ by the multiplication by a graph-to-norm continuous map ${{}^{g}\mathcal{R}}(H,H^{\prime})\to\mathcal{U}(H)$. Therefore, our construction determines $v$ up to the right multiplication by such a map.
Proof of Theorem B.   In order to prove the first part of the theorem we need to show that $\Phi$ is homotopic to the identity as a self-map of both ${{}^{r}\mathcal{R}}_{K}$ and ${{}^{g}\mathcal{R}}_{K}$. For the second part we need to prove a similar result for ${{}^{r}\mathcal{R}}_{F}$ and ${{}^{g}\mathcal{R}}_{F}$.
@@
In order to finish the proof of the Fredholm part, we need to show that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are homotopic. This part is more delicate because the action of ${{}^{s}{\mathcal{U}}}$ on ${{}^{g}\mathcal{R}}_{F}$ is not continuous, so we cannot directly apply a contraction of ${{}^{s}{\mathcal{U}}}$ for this purpose. However, we can use a part of [$P_{3}$, Theorem A] which states that the embedding ${{}^{g}\mathcal{R}}_{K}\hookrightarrow{{}^{g}\mathcal{R}}_{F}$ is a homotopy equivalence. We already proved that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ are homotopic; it follows that $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are also homotopic. Therefore, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$. $\square$
Proof of Theorem C.   Suppose that a global trivialization of $\mathcal{H}^{\prime}$ is fixed. Choose an arbitrary trivialization $\tau\colon\mathcal{H}\to X\times H$ of $\mathcal{H}$. In these trivializations $\mathcal{A}$ becomes a graph continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$. Let $v\colon{{}^{g}\mathcal{R}}(H,H^{\prime})\to{{}^{s}{\mathcal{U}}}(H)$ be a map provided by Theorem A. Then the map $X\to{{}^{s}{\mathcal{U}}}(H)$, $x\mapsto v(\mathcal{A}_{x})$ is continuous, so
@@
If instead a global trivialization of $\mathcal{H}$ is fixed, then one can interchange $\mathcal{H}$ and $\mathcal{H}^{\prime}$ by passing to the adjoint operators. Since the map $A\mapsto A^{*}$ is both graph and Riesz continuous, it follows from the first part of the proof that there is a global trivialization of $\mathcal{H}^{\prime}$ taking $\mathcal{A}$ to a Riesz continuous map. $\square$
$K^{0}$ index for graph continuous families.   By a classical result of Atiyah and Jänich, the space $\mathcal{B}_{F}(H)$ of bounded Fredholm operators is a classifying space for the functor $K^{0}$. See [A], Theorem A1. For an arbitrary parameter space $X$, $K^{0}(X)$ might be defined as the set $[X,\mathcal{B}_{F}]$ of homotopy classes of maps $X\to\mathcal{B}_{F}(H)$, with the group structure on $[X,\mathcal{B}_{F}]$ introduced in the standard way. The class of a continuous map $\alpha\colon X\to\mathcal{B}_{F}$ in $K^{0}(X)$ is called the index of $\alpha$:
@@
It can be easily seen, using linear homotopies, that the bounded transform map $\bm{\chi}\colon{{}^{r}\mathcal{R}}_{F}\to\mathcal{B}_{F}$ is homotopy inverse to the embedding $\mathcal{B}_{F}\hookrightarrow{{}^{r}\mathcal{R}}_{F}$, so both these maps are homotopy equivalencies. Therefore, the index of a Riesz continuous family of Fredholm operators is naturally defined as the index of its bounded transform:
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{r}\mathcal{R}}_{F}.$
Let now $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ be a graph continuous family of Fredholm operators. The space ${{}^{g}\mathcal{R}}_{F}$ is also a classifying space for the functor $K^{0}$, see [J], Theorem 3.5(i) or [$P_{3}$], Theorem A. Hence the homotopy class of $\mathcal{A}$ determines an element of $K^{0}(X)$ which can be called the index of $\mathcal{A}$:
@@
By the definition, two graph continuous families of Fredholm operators are homotopic if and only if their indices coincide.
@@
Theorems A and B of the present paper allow to relate the index of $\mathcal{A}$ given by (3.1) to the Fredholm index of bounded operators as follows. The composition of $\mathcal{A}$ with a map $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ provided by Theorem A is Riesz continuous, so $\bm{\chi}\circ\Phi\circ\mathcal{A}$ is a norm continuous family of bounded Fredholm operators. Theorem B provides the equality
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\Phi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F},$
@@
which can be used as the definition of the index of $\mathcal{A}$. Note that Theorem B implies that the homotopy class of $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{r}\mathcal{R}}_{F}$ does not depend on the choice of $v$, so $\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$ is also independent of this choice. Equivalently, the index can be defined in terms of graphs of operators as described in the Introduction.
For a Riesz continuous family $\mathcal{A}\colon X\to\mathcal{R}_{F}$ (in particular, for a norm continuous family of bounded Fredholm operators) the two definitions of the index coincide. Indeed, $\Phi(\mathcal{A}_{x})=\mathcal{A}_{x}\cdot v(\mathcal{A}_{x})$, so $\Phi\circ\mathcal{A}$ differs from $\mathcal{A}$ by the multiplication by a map $v\circ\mathcal{A}\colon X\to\mathcal{U}(H)$. If $\mathcal{A}$ is Riesz continuous, then $v\circ\mathcal{A}$ is norm continuous. Since $\mathcal{U}(H)$ is contractible, $\Phi\circ\mathcal{A}$ and $\mathcal{A}$ are homotopic as maps from $X$ to ${{}^{r}\mathcal{R}}_{F}$ and thus their indices coincide.
@@
Recently N. Ivanov [$I_{1}$, $I_{2}$] developed a new approach to the analytical index of families. He introduced the notion of a “Fredholm family” and defined the $K^{0}(X)$-valued index, which we will denote by $\operatorname{Ind}$, for Fredholm families parametrized by points of a paracompact space $X$. See [$I_{2}$], Section 3. A Fredholm family is a family of Fredholm operators $\mathcal{A}_{x}$ such that the finite-dimensional subspaces $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$ and $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}^{*}_{x}|)$, as well as the restriction of $\mathcal{A}_{x}$ to $\operatorname{\mathbb{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$, depend continuously on $x$ for an appropriate locally chosen $\varepsilon>0$, without any additional continuity assumptions. Ivanov’s index $\operatorname{Ind}$ is defined both for operators acting on a fixed Hilbert space and on fibers of a Hilbert bundle; it is invariant under isomorphisms of Hilbert bundles and under passing to the bounded transform.
Let $\mathcal{A}$ be a graph continuous family of Fredholm operators in $H$ parametrized by points of a paracompact space $X$. Then $\mathcal{A}$ is a Fredholm family in Ivanov’s sense, and we claim that our definition of the index of $\mathcal{A}$ provides the same element of $K^{0}(X)$ as Ivanov’s definition. Indeed, the properties of $\operatorname{Ind}$ mentioned above imply the equality $\operatorname{Ind}(\mathcal{A})=\operatorname{Ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$. Consider the norm continuous family $\alpha=\bm{\chi}\circ\Phi\circ\mathcal{A}$ of bounded Fredholm operators in $H$. By Theorem 7.2 of [$I_{2}$], $\operatorname{Ind}\alpha$ coincides with the index of $\alpha$ defined by Segal in [Se]. The last index is equal to the homotopy class $[\alpha]$ of $\alpha$ in $[X,\mathcal{B}_{F}(H)]=K^{0}(X)$ by [Se], Theorem 5.1(b). Taking all this together, we obtain
$\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\alpha)=[\alpha]=\operatorname{Ind}(\alpha)=\operatorname{Ind}(\mathcal{A})\in K^{0}(X).$
4 Polarizations

### v3

@@ 2 Preliminaries
In fact, this reasoning provides a stronger result. It concerns another natural topology on $\mathcal{R}(H,H^{\prime})$ induced by the injection $\mathbf{p}\colon\mathcal{R}(H,H^{\prime})\hookrightarrow{{}^{s}{\mathcal{P}}}(H\oplus H^{\prime})$ from the strong operator topology on the space of projections. On $\mathcal{R}^{\mathrm{sa}}(H)$ it coincides with the strong resolvent topology. The reasoning in the proof of Proposition 2.1 shows that the bounded transform $\bm{\chi}\colon\mathcal{R}\to\mathcal{D}$ is continuous with respect to this coarser topology on $\mathcal{R}$ and the strong* operator topology on $\mathcal{D}$.
For an arbitrary function $f\in C_{b}(\mathbb{R})$, a similar reasoning provides continuity of the map $f\colon\mathcal{R}^{\mathrm{sa}}(H)\to\mathcal{B}(H)$ with respect to the strong resolvent topology on $\mathcal{R}^{\mathrm{sa}}$ and the strong topology on $\mathcal{B}$. A different proof of such continuity of $f$ based on the spectral theorem is given by [RS], Theorem VIII.20(b). Conversely, one can deduce Proposition 2.1 from that theorem taking $f(x)=x(1+x^{2})^{-1/2}$ and using the standard trick of passing from $A$ to the self-adjoint operator $\left(\begin{smallmatrix}0&A^{*}\\ A&0\end{smallmatrix}\right)$. However, the direct proof of the proposition given above and not based on the spectral theorem seems to be more transparent.
Action of the unitary groups.  The unitary groups $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ act on $\mathcal{R}(H,H^{\prime})$ by the right and left multiplication respectively. These actions preserve the subsets $\mathcal{R}_{F}$ and $\mathcal{R}_{K}$.
Proposition 2.2.
The actions of the groups $\mathcal{U}(H)$ and $\mathcal{U}(H^{\prime})$ (equipped with the norm topology) on ${{}^{r}\mathcal{R}}(H,H^{\prime})$ and ${{}^{g}\mathcal{R}}(H,H^{\prime})$ are continuous.
Proof. For $u\in\mathcal{U}(H^{\prime})$ and $v\in\mathcal{U}(H)$ we have
@@
which proves the proposition. $\square$
Let ${{}^{s}{\mathcal{U}}}(H)$ denote the unitary group equipped with the strong operator topology. It is a topological group, see [RW, Lemma 1.3]. The actions of ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on $^{r}\mathcal{R}$} and $^{g}\mathcal{R}$} are not continuous, but continuity of the action of $^{s}{\mathcal{U}}$} on $\mathcal{K}$ provides the following result.
@@
The actions of the groups ${{}^{s}{\mathcal{U}}}(H)$ and ${{}^{s}{\mathcal{U}}}(H^{\prime})$ on ${{}^{g}\mathcal{R}}_{K}(H,H^{\prime})$ are continuous.
@@
(2.8) $\mathcal{G}r_{K}(H\oplus H^{\prime})\cong\mathcal{P}_{K}(H\oplus H^{\prime})=\left\{p\in\mathcal{P}(H\oplus H^{\prime})\mid p-p_{\infty}\text{ is compact}\right\},$
@@
3 Proof of Theorem 1
The idea of the proof can be breifly described as follows. The graphs $\Gamma_{A}$ of regular operators $A\colon H\to H^{\prime}$ form fibers of a locally trivial Hilbert bundle $\Gamma$ over ${{}^{g}\mathcal{R}}(H,H^{\prime})$ with the structure group $\mathcal{U}(H)$. Since $\mathcal{U}(H)$ is contractible, this bundle admits a trivialization $u\colon{{}^{g}\mathcal{R}}\times H\stackrel{{\scriptstyle{\approx}\hskip 1.63885pt}}{{\longrightarrow}}\Gamma$. The composition of the unitary operator $u_{A}\colon H\stackrel{{\scriptstyle{\approx}\hskip 1.63885pt}}{{\longrightarrow}}\Gamma_{A}\subset H\oplus H^{\prime}$ with the projection $H\oplus H^{\prime}\to H^{\prime}$ provides the contraction $\varphi_{A}\colon H\to H^{\prime}$. We define $\Phi$ by the formula $\bm{\chi}(\Phi_{A})=\varphi_{A}$ and show that this map satisfies all conclusions of the theorem.
@@ 3 Proof of Theorem 1
$\mathcal{I}=\left\{u\in\mathcal{B}(H,H\oplus H^{\prime})\mid u^{*}u=\operatorname{Id}_{H}\right\}$
@@
$\bm{\pi}\colon\mathcal{I}\to\mathcal{G}r(H\oplus H^{\prime}),\quad u\mapsto\operatorname{Ran}(u).$
takes an isometry $u$ to its range; the corresponding projection is $uu^{*}$. The image $\bm{\pi}(\mathcal{I})$ of this map consists of infinite-dimensional subspaces of $H\oplus H^{\prime}$. The group $\mathcal{U}(H)$ acts on $\mathcal{I}$ by the right multiplication. Kuiper’s theorem [Ku] together with the same standard arguments as were used in the beginning of Section 2 show that $\bm{\pi}\colon\mathcal{I}\to\bm{\pi}(\mathcal{I})$ is a principal $\mathcal{U}(H)$-bundle admitting a continuous section (and thus is trivial).
Proof of Theorem 1.  The graphs of regular operators have infinite dimension, so ${{}^{g}\mathcal{R}}\subset\bm{\pi}(\mathcal{I})\subset\mathcal{G}r(H\oplus H^{\prime})$. For the rest of the proof we fix a section of $\bm{\pi}\colon\mathcal{I}\to\bm{\pi}(\mathcal{I})$ and denote by $u\colon{{}^{g}\mathcal{R}}\to\mathcal{I}$ its restriction to $^{g}\mathcal{R}$}. Then $u$ is a graph-to-norm continuous map taking a regular operator $A$ to an isometry $u_{A}\colon H\to H\oplus H^{\prime}$ with the range $\Gamma_{A}$.
@@
4 Applications of Theorem 1
Proof of Theorem 2.  In order to prove the first part of the theorem, we need to show that $\Phi$ is homotopic to the identity as a self-map of both ${{}^{r}\mathcal{R}}_{K}$ and ${{}^{g}\mathcal{R}}_{K}$. For the second part we need to prove a similar result for ${{}^{r}\mathcal{R}}_{F}$ and ${{}^{g}\mathcal{R}}_{F}$.
@@ 4 Applications of Theorem 1
In order to finish the proof of the Fredholm part, we need to show that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are homotopic. This part is more delicate because the action of $^{s}{\mathcal{U}}$} on ${{}^{g}\mathcal{R}}_{F}$ is not continuous, so we cannot directly apply a contraction of $^{s}{\mathcal{U}}$} for this purpose. However, we can use a part of [P3, Theorem A] which states that the embedding ${{}^{g}\mathcal{R}}_{K}\hookrightarrow{{}^{g}\mathcal{R}}_{F}$ is a homotopy equivalence. We already proved that the maps $\Phi\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{K}\to{{}^{g}\mathcal{R}}_{K}$ are homotopic; it follows that $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ and $\operatorname{Id}\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$ are also homotopic. Therefore, the restriction of $\Phi$ to the subspace of Fredholm operators is homotopy inverse to the identity map ${{}^{r}\mathcal{R}}_{F}\to{{}^{g}\mathcal{R}}_{F}$. $\square$
Proof of Theorem 3.  Suppose that a global trivialization of $\mathcal{H}^{\prime}$ is fixed. Choose an arbitrary trivialization $\tau\colon\mathcal{H}\to X\times H$ of $\mathcal{H}$. In these trivializations $\mathcal{A}$ becomes a graph continuous map $X\to\mathcal{R}_{K}(H,H^{\prime})$. Let $v\colon{{}^{g}\mathcal{R}}(H,H^{\prime})\to{{}^{s}{\mathcal{U}}}(H)$ be a map provided by Theorem 1. Then the map $X\to{{}^{s}{\mathcal{U}}}(H)$, $x\mapsto v(\mathcal{A}_{x})$ is continuous, so
@@
If instead a global trivialization of $\mathcal{H}$ is fixed, then one can pass to the adjoint operators, swapping $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Since the map $A\mapsto A^{*}$ is both graph and Riesz continuous, it follows from the first part of the proof that there is a global trivialization of $\mathcal{H}^{\prime}$ taking $\mathcal{A}$ to a Riesz continuous map. $\square$
$K^{0}$ index for graph continuous families.  By a classical result of Atiyah and Jänich, the space $\mathcal{B}_{F}(H)$ of bounded Fredholm operators is a classifying space for the functor $K^{0}$. See [A], Theorem A1. For an arbitrary parameter space $X$, $K^{0}(X)$ might be defined as the set $X,\mathcal{B}_{F}]$ of homotopy classes of maps $X\to\mathcal{B}_{F}(H)$, with the group structure on $X,\mathcal{B}_{F}]$ introduced in the standard way. The class of a continuous map $\alpha\colon X\to\mathcal{B}_{F}$ in $K^{0}(X)$ is called the index of $\alpha$:
@@
It can be easily seen, using linear homotopies, that the bounded transform $\bm{\chi}\colon{{}^{r}\mathcal{R}}_{F}\to\mathcal{B}_{F}$, from regular Fredholm operators to bounded Fredholm operators, is homotopy inverse to the embedding $\mathcal{B}_{F}\hookrightarrow{{}^{r}\mathcal{R}}_{F}$, so both these maps are homotopy equivalencies. Therefore, the index of a Riesz continuous map $\mathcal{A}\colon X\to{{}^{r}\mathcal{R}}_{F}$ is naturally defined as the index of its bounded transform:
(4.1) $\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{r}\mathcal{R}}_{F}.$
Let now $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ be a graph continuous map. The space ${{}^{g}\mathcal{R}}_{F}$ is also a classifying space for the functor $K^{0}$, see [J], Theorem 3.5(i) or [P3], Theorem A. Hence the homotopy class of $\mathcal{A}$ determines an element of $K^{0}(X)$ which can be called the index of $\mathcal{A}$:
@@
Theorems 1 and 2 of the present paper allow to relate the index of $\mathcal{A}$ given by (4.2) to the Fredholm index of bounded operators as follows. The composition of $\mathcal{A}$ with a map $\Phi\colon{{}^{g}\mathcal{R}}\to{{}^{r}\mathcal{R}}$ provided by Theorem 1 is Riesz continuous, so $\bm{\chi}\circ\Phi\circ\mathcal{A}$ is a norm continuous map $X\to\mathcal{B}_{F}$. Theorem 2 implies
(4.3) $\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\Phi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)\;\;\text{for}\;\;\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F},$
@@
which can be used as the definition of the index of $\mathcal{A}$. It follows from Theorem 2 that the homotopy class of $\Phi\colon{{}^{g}\mathcal{R}}_{F}\to{{}^{r}\mathcal{R}}_{F}$ is independent of the choice of $v$, so $\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$ is also independent of this choice.
Similarly, Theorem 2 implies that for a Riesz continuous map $\mathcal{A}\colon X\to\mathcal{R}_{F}$ (in particular, for a norm continuous map $X\to\mathcal{B}_{F}$) the two definitions of the index given by (4.1) and (4.2) coincide.
@@
Recently N. Ivanov [I1, I2] developed a new approach to the family index. He introduced the notion of a “Fredholm family” and defined a $K^{0}(X)$-valued index, which we will denote by $\operatorname{Ind}$, for Fredholm families parametrized by points of a paracompact space $X$. See [I2], Section 3. A Fredholm family is a family of Fredholm operators $\mathcal{A}_{x}$ such that the finite-dimensional subspaces $\operatorname{\mathds{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$ and $\operatorname{\mathds{1}}_{[0,\varepsilon]}(|\mathcal{A}^{*}_{x}|)$, as well as the restriction of $\mathcal{A}_{x}$ to $\operatorname{\mathds{1}}_{[0,\varepsilon]}(|\mathcal{A}_{x}|)$, depend continuously on $x$ for an appropriate locally chosen $\varepsilon>0$, without any additional continuity assumptions. Ivanov’s index $\operatorname{Ind}$ is defined both for operators acting on a fixed Hilbert space and on fibers of a Hilbert bundle; it is invariant under isomorphisms of Hilbert bundles and under passing to the bounded transform. By the construction, for a norm continuous map $\mathcal{A}\colon X\to\mathcal{B}_{F}$ Ivanov’s index coincides with the usual index $\operatorname{ind}(\mathcal{A})\in[X,\mathcal{B}_{F}]=K^{0}(X)$.
Let now $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ be a graph continuous map from a paracompact space $X$. Then $\mathcal{A}$ is a Fredholm family in Ivanov’s sense, and the properties of $\operatorname{Ind}$ mentioned above imply the equality
$\operatorname{Ind}(\mathcal{A})=\operatorname{Ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A}).$
Therefore, our definition of the index of $\mathcal{A}$ provides the same element of $K^{0}(X)$ as Ivanov’s definition.
Remark.
Our proof of Theorem 1 shows that the definition of the index of $\mathcal{A}\colon X\to{{}^{g}\mathcal{R}}_{F}$ given by the formula $\operatorname{ind}(\mathcal{A})=\operatorname{ind}(\bm{\chi}\circ\Phi\circ\mathcal{A})$ can be equivalently stated in terms of the graphs $\Gamma_{x}$ of operators $\mathcal{A}_{x}$, as the index of the family of projections $\Gamma_{x}\to H^{\prime}$. For this, one needs to trivialize the Hilbert bundle $\Gamma$, formed by subspaces $\Gamma_{x}\subset H\oplus H^{\prime}$, in the category of bundles with the structure group $\mathcal{U}$. Note that such a trivialization is not canonical; it is defined only up to a continuous map $X\to\mathcal{U}(H)$. However, change of a trivialization does not affect the homotopy type of the corresponding family of bounded Fredholm operators, since the unitary group $\mathcal{U}(H)$ is contractible.
5 Polarizations and proof of Theorem 4